OpenStudy (unklerhaukus):

Find the inverse laplace transform of the arctangent of 1/p \[\mathcal L^{-1}\{\arctan\tfrac1p\}\]

OpenStudy (anonymous):

You have to find it?

OpenStudy (unklerhaukus):

yeah, i dont know how to start

OpenStudy (anonymous):

I think working with the function's derivative might help. \[\frac{d}{dp}\left[\arctan\frac{1}{p}\right]=-\frac{1}{p^2+1}\]

OpenStudy (anonymous):

So in effect, you'd be finding the inverse transform of an integral of this function. Are there any rules you know about the transform of an integral?

OpenStudy (unklerhaukus):

something like this

OpenStudy (anonymous):

sin(t)/t

OpenStudy (unklerhaukus):

\[\mathcal L^{-1}\{\arctan\tfrac1p\}=x(t)\\ -\frac1{(1/p)^2+1}={X(p)}\\ -\frac{p^2}{1+p^2}={X(p)}\\ \] am i on the right track?

OpenStudy (anonymous):

If \(\dfrac{\sin t}{t}\) is the answer, then using the rule is the right way to go.

OpenStudy (anonymous):

\[L(\sin(t)=1/1+s2=d(\tan(p))/dp L(\sin(t)/t=\int\limits_{p}^{infinity}1/1+s2ds=\pi/2-\tan-1s=\cot-1s=\tan-1(1/s)\]

OpenStudy (unklerhaukus):

\[\newcommand \p \newcommand \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \p \Lin [1] {\operatorname{\mathcal L}^{-1}\left\{#1\right\} } \p \intl [4] {\int\limits_{#1}^{#2}{#3}\dd{#4} } \begin{align*} \Lin{\arctan\tfrac1p} &=\Lin{\intl p\infty{\de{}p(\arctan\tfrac1p)} p} \\ &=\Lin{\intl p\infty{\frac{-1/p^2}{1+(1/p)^2}}p} \\ &=\Lin{\intl p\infty{\frac{-1}{p^2+1}}p} \\ &=\frac{\sin t}t \\ \end{align*}\]

OpenStudy (unklerhaukus):

is this right ?

OpenStudy (usukidoll):

differential equations...I'm not even there yet...

OpenStudy (unklerhaukus):

i'm still not understanding this completely

OpenStudy (unklerhaukus):

Hows this? \[ \newcommand \p \newcommand \p \Lap [1] {\operatorname{\mathcal L}\left\{\rule{0pt}{2.2ex}#1\right\} } \p \dd [1] {\,\mathrm d#1 } \p \de [2] {\frac{\mathrm d #1}{\mathrm d#2} } \p \Lin [1] {\operatorname{\mathcal L}^{-1}\left\{#1\right\} } \p \intl [4] {\int\limits_{#1}^{#2}{#3}\dd{#4} } \begin{align*} \Lin{\arctan\tfrac1p} &=\Lin{\intl p\infty{F(p)}p} =\frac{f(t)}t\\ \\ \arctan\tfrac1p &=\intl p\infty{F(p)}p \\ \de{}p\left(\arctan\tfrac1p\right) &=\de{}p\intl p\infty{F(p)}p \\ \frac1{1+1/p^2}\cdot-1/p^2 &=F(\infty)\cdot(\infty)'-F(p)\cdot(p)' \\ \frac{-1/p^2}{1+1/p^2} &=-F(p) \\ \frac{1}{p^2+1} &=F(p) \\ \\ \Lin{\frac{1}{p^2+1}} &=\Lin{F(p)} \\ \sin(t) &=f(t) \\ \\ \Lin{\arctan\tfrac1p} &=\frac{\sin t}t \end{align*} \]

OpenStudy (anonymous):

I'm not sure I do either, my class and textbook didn't cover Laplace past basic inverse transforms.